OpenStudy (anonymous):

Determine the lim ((1/(x-1))-(1/(x^2-x)) x->1

OpenStudy (anonymous):

\[\frac{1}{x-1}-\frac{1}{x^2-x}=\frac{x-1}{x^2-x}=\frac{x-1}{x(x-1)}=\frac{1}{x}\]

OpenStudy (anonymous):

replace x by 1 to get your answer